# Physics-informed approach for exploratory Hamilton–Jacobi–Bellman equations via policy iterations

### v1

@@ 1 Introduction
Solving nonlinear Hamilton–Jacobi–Bellman (HJB) equations lies at the core of stochastic optimal control. In continuous-time settings, control strategies must handle uncertainty while ensuring long-term optimality. Among various formulations, entropy-regularized control which is also known as exploratory or soft control has emerged as a powerful paradigm that augments the running cost with a Kullback–Leibler divergence penalty against a reference measure. This regularization not only induces stochasticity in control policies but also promotes robust exploration and analytical tractability (Wang, Zariphopoulou, and Zhou 2020; Ziebart et al. 2008). While entropy-regularized control in continuous time (Wang, Zariphopoulou, and Zhou 2020) is rigorously formulated, the work does not address the challenges of function approximation or high-dimensional computation. Our framework complements this theory by incorporating neural approximation and residual-driven learning in a mesh-free manner.
The resulting HJB equation is a linear second-order elliptic partial differential equation (PDE), typically defined over an unbounded domain. Solving such equations numerically remains a central challenge, especially in high dimensions. Soft policy iteration (PI) (Tran, Wang, and Zhang 2025; Tang, Zhang, and Zhou 2022) alternates between policy improvement via a soft-max update and policy evaluation via solving a linear elliptic PDE. Their analysis rigorously establishes exponential convergence under two canonical regimes: bounded coefficients with small control in the diffusion, and unbounded dynamics under sufficient discounting. Their results extend classical Howard-type PI into the entropy-regularized setting with relaxed controls and randomized policies.
While the theory of soft PI is well understood, practical deployment in high-dimensional control remains difficult because each iteration requires solving a nonlinear PDE. Recent work has attempted to bypass grids with mesh-free physics-informed neural networks (PINNs), embedding the PDE residual into a variational loss (Lee and Kim 2025; Ramesh and Ravindran 2023). Yet rigorous a priori guarantees in the entropy-regularized setting are still missing. In the deterministic case, PINN-based policy iteration methods (Meng et al. 2024; Lee and Kim 2025) establish only $L^{\infty}$ convergence on compact domains and do not incorporate entropy regularization or soft-max policies. Separately, a physics-informed model-based reinforcement learning framework (Ramesh and Ravindran 2023) constrains the learned dynamics model but does not formulate or analyze the underlying HJB. In contrast, our method targets the entropy-regularized HJB directly and derives an $L^{2}$ error analysis for a fully mesh-free PI scheme.
@@ 2.1 Exploratory HJB equations
Let $U\subset\mathbb{R}^{m}$ denote the compact control space. We consider a filtered probability space $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ equipped with a $d$-dimensional Brownian motion $(W_{t})_{t\geq 0}$. For each relaxed control $\pi=\{\pi(x,\cdot)\}_{x\in\mathbb{R}^{d}}$, where $\pi(x,\cdot)\in\mathbb{P}(U)$ is a probability measure on $U$ for each $x$, the state evolves according to the following SDE
@@
×$\quad\ \ \ \times\left(\left.\int_{U}\left(r(X^{\pi}_{t},u)-\lambda\ln\pi(X_{t},u)\right)\pi(X_{t},u)du\right)dt\right].$
@@
[…]
@@ 2.2 Policy iteration scheme
In Algorithm 1, we present the exact policy iteration scheme proposed in the soft PI framework (Tang, Zhang, and Zhou 2022; Tran, Wang, and Zhang 2025), which we later extend to a mesh-free, PINN-based implementation. The algorithm alternates between two steps: a policy improvement step, where a soft-optimal policy is computed via the Boltzmann distribution, and a policy evaluation step, which involves solving a linear elliptic PDE under the fixed policy. The key advantage of this formulation is that the original HJB equation is decomposed into a sequence of tractable subproblems, specifically, linear PDEs with frozen coefficients and closed-form policy updates. This structure not only simplifies analysis but also lends itself naturally to function approximation methods.
We initialize the algorithm with a value function $v^{0}\in C_{b}^{2}(\mathbb{R}^{d})$, which may be chosen as a zero function or a smooth prior based on approximate linearization (e.g., a quadratic ansatz from LQR theory). While the initial guess affects the transient phase, the policy iteration scheme (Tran, Wang, and Zhang 2025; Tang, Zhang, and Zhou 2022) is contractive and converges exponentially to the optimal value regardless of $v^{0}$.
@@
For completeness, we collect the structural assumptions and key convergence results (Tran, Wang, and Zhang 2025; Ma, Wang, and Zhang 2024), as our $L^{2}$-analysis in Section 3.1 builds on these facts.
@@
$B:=\sup_{a\in U}\|\nabla_{x}\cdot b(\cdot,a)\|_{L^{\infty}(\mathbb{R}^{d})}<\infty$, and there exists $R>0$ such that $b(a,\cdot)=\sigma(\cdot)=0$ for $|x|\geq R$.
@@
Given $\tilde{r},\tilde{b},\tilde{\sigma}\in C_{b}^{2}(\mathbb{R}^{d})$, assume that (1) $\rho>\frac{1}{2}B$, where $B:=\|\nabla_{x}\cdot\tilde{b}(\cdot,a)\|_{L^{\infty}(\mathbb{R}^{d})}$, (2) $\tilde{\Sigma}(x)\succeq\frac{1}{C_{0}}I_{d}$, (3) $\tilde{r}\in L^{2}(\mathbb{R}^{d})$. If $v$ is a unique classical solution to
$\rho v(x)-\tilde{r}(x)-\tilde{b}(x)\cdot\nabla_{x}v(x)-\tfrac{1}{2}\mathrm{tr}(\tilde{\Sigma}(x)D_{xx}^{2}v(x))=0,$
@@
To control how approximation errors in the value function affect the policy, we analyze the Lipschitz continuity of the softmax map. The following result shows that the mapping from $\nabla_{x}v$ to the induced policy $\pi[v]$ is Lipschitz-continuous in the $L^{2}(U)$ norm, with an explicit constant.
@@
We fix a state $x\in\mathbb{R}^{d}$ and the control variable $u\in U\subset\mathbb{R}^{m}$. Under Assumption 1, and
@@
for all $|p|,|q|\leq\tilde{L}>0$. Here, the constant
$L_{\Phi}:=\frac{2\sup_{|p|\leq\tilde{L}}\Phi(p)\sqrt{d}|U|^{1/2}}{\lambda}.$ (7)
@@
Differentiating the softmax map $\Phi(p)$ with respect to $p$ reveals a Jacobian structure involving $\Phi(p)$ itself and the control vector $b(x,u)$. The gradient is uniformly bounded due to the boundedness of $b$, and integration over $U$ yields the global Lipschitz constant $L_{\Phi}$. A full derivation appears in Appendix 2.
Having established the regularity properties of the soft-optimal policy, we now recall the global convergence guarantee for the exact soft policy iteration scheme (Tran, Wang, and Zhang 2025; Tang, Zhang, and Zhou 2022). This result confirms that the sequence $\{v^{n}\}$ generated by Algorithm 1 converges exponentially fast to the unique solution $V$ of the entropy-regularized HJB equation. It forms the backbone of our error decomposition in Section 3.1, where we extend this result to the approximate PINN-based implementation.
@@ 3 Physics-informed mesh-free approach
8:Check convergence: If $\frac{1}{N}\sum_{i=1}^{N}|v(x_{i};\theta_{n+1})-v(x_{i};\theta_{n})|^{2}<\varepsilon$then stop.
@@ 3.1 Error estimates
$b(x,a)=0\quad\text{and}\quad\sigma(x)=0\quad\text{for all }|x|\geq R,\ a\in U.$
@@
$\tilde{v}^{n}-V=\underbrace{\tilde{v}^{n}-\hat{v}^{n}}_{\displaystyle\text{(PDE error)}}+\underbrace{\hat{v}^{n}-v^{n}}_{\displaystyle\text{(policy error)}}+\underbrace{v^{n}-V}_{\displaystyle\text{(iteration error)}}.$
@@
$\begin{split}&r_{n}(x):=\|\tilde{\pi}^{n}(x,\cdot)-\hat{\pi}^{n}(x,\cdot)\|_{L^{2}(U)},\\ &q_{n}(x):=\rho\tilde{v}^{n}-F_{\tilde{\pi}^{n}}(x,\nabla_{x}v^{n},D_{xx}^{2}\tilde{v}^{n})\end{split}$
@@
The quantity $\|r_{n}\|_{L^{2}(\mathcal{X})}$ and $\|q_{n}\|_{L^{2}(\mathcal{X})}$ measure the PDE residual loss and policy improvement loss, respectively. We note here that we will derive a convergence result in terms of $r_{n}$ rather than the KL-divergence between $\tilde{\pi}^{n}$ and $\hat{\pi}^{n}$ as we have that
@@
from the Pinsker’s inequality (Csiszár and Körner 2011).
@@ 4 Experiments
We evaluate the proposed PINN-based soft policy iteration (PINN-SPI) framework on a suite of entropy-regularized stochastic control problems, ranging from low-dimensional nonlinear systems to high-dimensional linear-quadratic regulators (LQR). Our goals are to (i) demonstrate scalability to high-dimensional settings, (ii) demonstrate the monotonicity property (Tran, Wang, and Zhang 2025)[Corollary 5.1].
@@ 4.1 Linear-Quadratic Regulator (LQR) with Compact Action Space
We consider entropy-regularized linear-quadratic regulator (LQR) problems in 5 and 10 dimensions with compact action constraints. The system dynamics are
@@
$\in U:=\{u\in\mathbb{R}^{m}\mid\|u\|_{\infty}\leq\textbf{u}\}.$ u } .
@@
We apply our PINN-based soft policy iteration (PINN-SPI) method and compare it to the model-free Soft Actor-Critic (SAC) algorithm. Both methods are initialized with the same linear controller derived from the unconstrained problem. Our approach uses residual minimization with randomly sampled collocation points and softmax-based policy updates.
@@
[…]
@@ Results.
Figure 1 compares PINN-SPI against Soft Actor-Critic (SAC) in 5D and 10D LQR settings. PINN-SPI consistently achieves higher reward and smoother convergence, while SAC struggles to generalize in high dimensions due to sample inefficiency.
Figure 2 further illustrates the convergence behavior of PINN-SPI. The evaluation reward increases monotonically over training time, confirming the theoretical stability of policy iteration. The 10D case in particular highlights the scalability of our mesh-free implementation.
@@
Figure 1: Comparison between PINN-SPI (ours) and SAC in learning stochastic entropy-regularized LQR problems.
(a) 5D LQR: PINN-SPI steadily improves total reward.
(b) 10D LQR: Stable and monotonic convergence.
@@ 4.2 Nonlinear Benchmarks with Stochastic Dynamics
To demonstrate performance in more realistic and nonlinear settings, we evaluate our method on two standard control benchmarks: the stochastic inverted pendulum and cartpole. Each system is modeled as a stochastic control-affine system with additive Brownian noise.
We adopt a standard stochasticized version of the cartpole and pendulum dynamics supported by OpenAI GYM (Brockman et al. 2016) with additive noise on state variables.
@@
For each environment, we define entropy-regularized cost functionals and apply PINN-SPI with a fixed noise level $\sigma=0.1\cdot I_{d}$. We use neural networks for both value approximation and policy extraction, trained via policy iteration with residual loss minimization. SAC is used as the primary model-free baseline, trained under the same noise realizations and reward structures.
@@
Figure 3 shows the performance over elapsed time. PINN-SPI achieves faster stabilization and higher reward than SAC, while maintaining constraint satisfaction. Especially under high noise, SAC tends to exhibit oscillatory behavior due to reward shaping issues, whereas PINN-SPI produces smoother trajectories guided by the physics-informed value landscape.
(a) Cartpole
(b) Pendulum
Figure 3: Comparison between PINN-SPI and SAC

### v2

@@ 1 Introduction
Solving nonlinear Hamilton–Jacobi–Bellman (HJB) equations lies at the core of stochastic optimal control. In continuous-time settings, control strategies must handle uncertainty while ensuring long-term optimality. Among various formulations, entropy-regularized control which is also known as exploratory or soft control has emerged as a powerful paradigm that augments the running cost with a Kullback–Leibler divergence penalty against a reference measure. This regularization not only induces stochasticity in control policies but also promotes robust exploration and analytical tractability (wang2020reinforcement; ziebart2008maximum). While entropy-regularized control in continuous time (wang2020reinforcement) is rigorously formulated, the work does not address the challenges of function approximation or high-dimensional computation. Our framework complements this theory by incorporating neural approximation and residual-driven learning in a mesh-free manner.
The resulting HJB equation is a nonlinear second-order elliptic partial differential equation (PDE), typically defined over an unbounded domain. Solving such equations numerically remains a central challenge, especially in high dimensions. Soft policy iteration (PI) (tran2025policy; tang2022exploratory) alternates between policy improvement via a soft-max update and policy evaluation via solving a linear elliptic PDE. Their analysis rigorously establishes exponential convergence under two canonical regimes: bounded coefficients with small control in the diffusion, and unbounded dynamics under sufficient discounting. Their results extend classical Howard-type PI into the entropy-regularized setting with relaxed controls and randomized policies.
While the theory of soft PI is well understood, practical deployment in high-dimensional control remains difficult because each iteration requires solving a nonlinear PDE. Recent work has attempted to bypass grids with mesh-free physics-informed neural networks (PINNs), embedding the PDE residual into a variational loss (lee2025hamilton; ramesh2023physics). Yet rigorous a priori guarantees in the entropy-regularized setting are still missing. In the deterministic case, PINN-based policy iteration methods (meng2024physics; lee2025hamilton) establish only $L^{\infty}$ convergence on compact domains and do not incorporate entropy regularization or soft-max policies. Separately, a physics-informed model-based reinforcement learning framework (ramesh2023physics) constrains the learned dynamics model but does not formulate or analyze the underlying HJB. In contrast, our method targets the entropy-regularized HJB directly and derives an $L^{2}$ error analysis for a fully mesh-free PI scheme.
@@ 2.1 Exploratory HJB equations
Let $U\subset\mathbb{R}^{m}$ denote the compact control space. We consider a filtered probability space $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ equipped with a $d$-dimensional Brownian motion $(W_{t})_{t\geq 0}$. For each relaxed control $\pi=\{\pi(x,\cdot)\}_{x\in\mathbb{R}^{d}}$, where $\pi(x,\cdot)\in\mathbb{P}(U)$ is a probability measure on $U$ for each xx, [^1] We consider relaxed controls $\pi(x,\cdot)\in\mathbb{P}(U)$ that admit densities with respect to the Lebesgue measure on $U$, so that $\pi(x,\,\mathrm{d}u)=\pi(x,u)\,\mathrm{d}u$ and all integrals $\int_{U}(\cdot)\pi(x,u)\,du$ are well-defined. the state evolves according to the following SDE
@@
$\quad\ \ \ \times\left(\left.\int_{U}\left(r(X^{\pi}_{t},u)-\lambda\ln\pi(X_{t},u)\right)\pi(X_{t},u)\,\mathrm{d}u\right)\,\mathrm{d}t\right].$
@@
Let us begin with some notations used throughout the paper. For $x\in\mathbb{R}^{d}$, we write $|x|$ for the Euclidean norm. We denote the Hessian of a scalar function $f$ by $D_{xx}^{2}f$. For $p\in[1,\infty)$, we write $f\in L^{p}(\Omega)$, if
@@ 2.2 Policy iteration scheme
In Algorithm 1, we present the exact policy iteration scheme proposed in the soft PI framework (tang2022exploratory; tran2025policy), which we later extend to a mesh-free, PINN-based implementation. The algorithm alternates between two steps: a policy improvement step, where a soft-optimal policy is computed via the Boltzmann distribution, and a policy evaluation step, which involves solving a linear elliptic PDE under the fixed policy. The key advantage of this formulation is that the original HJB equation is decomposed into a sequence of tractable subproblems, specifically, linear PDEs with frozen coefficients and closed-form policy updates. This structure not only simplifies analysis but also lends itself naturally to function approximation methods.
We initialize the algorithm with a value function $v^{0}\in C_{b}^{2}(\mathbb{R}^{d})$, which may be chosen as a zero function or a smooth prior based on approximate linearization (e.g., a quadratic ansatz from LQR theory). While the initial guess affects the transient phase, the policy iteration scheme (tran2025policy; tang2022exploratory) is contractive and converges exponentially to the optimal value regardless of $v^{0}$.
@@
For completeness, we collect the structural assumptions and key convergence results (tran2025policy; ma2024convergence), as our $L^{2}$-analysis in Section 3.1 builds on these facts.
@@
$B:=\sup_{u\in U}\|\nabla_{x}\cdot b(\cdot,u)\|_{L^{\infty}(\mathbb{R}^{d})}<\infty$, and there exists $R>0$ such that $b(x,u)=0$ uniformly in $u\in U$ for $|x|\geq R$.
@@
Given $\tilde{r},\tilde{b},\tilde{\sigma}\in C_{b}^{2}(\mathbb{R}^{d})$, assume that (i) $\rho>\frac{1}{2}B$, where $B:=\|\nabla_{x}\cdot\tilde{b}(\cdot,u)\|_{L^{\infty}(\mathbb{R}^{d})}$, (ii) $\tilde{\Sigma}(x)\succeq\frac{1}{C_{0}}I_{d}$, (iii) $\tilde{r}\in L^{2}(\mathbb{R}^{d})$. Let $v$ be a unique classical solution to
$\rho v(x)-\tilde{r}(x)-\tilde{b}(x)\cdot\nabla_{x}v(x)-\tfrac{1}{2}\mathrm{tr}(\tilde{\Sigma}(x)D_{xx}^{2}v(x))=0.$
@@
To control how approximation errors in the value function affect the policy, we analyze the Lipschitz continuity of the softmax map. The following result shows that the mapping from $\nabla_{x}v$ to the induced policy $\pi[v]$ is Lipschitz continuous in the $L^{2}(U)$ norm, with an explicit constant.
@@
We fix a state $x\in\mathbb{R}^{d}$ and the control variable $u\in U\subset\mathbb{R}^{m}$. Under Assumption 1, we have that
@@
for all $|p|,|q|\leq\tilde{L}$. Here, the constant
$L_{\Phi}:=L_{\phi}:=\frac{2\|b\|_{L^{\infty}}\sup_{|p|\leq\tilde{L}}\Phi(p)\sqrt{d}|U|^{1/2}}{\lambda}.$ (7)
@@
Differentiating the softmax map $\Phi(p)$ with respect to $p$ reveals a Jacobian structure involving $\Phi(p)$ itself and the control vector $b(x,u)$. The gradient is uniformly bounded due to the boundedness of $b$, and integration over $U$ yields the global Lipschitz constant $L_{\Phi}$. A full derivation appears in Appendix B.
Having established the regularity properties of the soft-optimal policy, we now recall the global convergence guarantee for the exact soft policy iteration scheme (tran2025policy; tang2022exploratory). This result confirms that the sequence $\{v^{n}\}$ generated by Algorithm 1 converges exponentially fast to the unique solution $V$ of the entropy-regularized HJB equation. It forms the backbone of our error decomposition in Section 3.1, where we extend this result to the approximate PINN-based implementation.
@@ 3 Physics-informed mesh-free approach
[…]
to minimize the KL divergence between $\hat{\pi}$ and $\pi$, where
@@ 3.1 Error estimates
$b(x,u)=0\quad\quad\text{for all }|x|\geq R,\quad u\in U.$
@@
$\tilde{v}^{n}-V=\underbrace{\tilde{v}^{n}-\hat{v}^{n}}_{\displaystyle\text{(PDE error)}}+\underbrace{\hat{v}^{n}-v^{n}}_{\displaystyle\text{(policy error)}}+\underbrace{v^{n}-V.}_{\displaystyle\text{(iteration error)}}$
@@
$\begin{split}&r_{n}(x):=\|\tilde{\pi}^{n}(x,\cdot)-\hat{\pi}^{n}(x,\cdot)\|_{L^{2}(U)},\\ &q_{n}(x):=\rho\tilde{v}^{n}-F_{\tilde{\pi}^{n}}(x,\nabla_{x}\tilde{v}^{n},D_{xx}^{2}\tilde{v}^{n}),\end{split}$
@@
The quantity $\|r_{n}\|_{L^{2}(\mathcal{X})}$ and $\|q_{n}\|_{L^{2}(\mathcal{X})}$ measure the policy improvement loss and PDE residual loss, respectively. We note here that we will derive a convergence result in terms of $r_{n}$ rather than the KL-divergence between $\tilde{\pi}^{n}$ and $\hat{\pi}^{n}$ as we have that
@@
from the Pinsker’s inequality (csiszar2011information).
@@ 4 Experiments
We evaluate the proposed PINN-based soft policy iteration (PINN-SPI) framework on a suite of entropy-regularized stochastic control problems, ranging from low-dimensional nonlinear systems to high-dimensional linear-quadratic regulators (LQR). Our goals are to (i) demonstrate scalability to high-dimensional settings, (ii) demonstrate the monotonicity property (tran2025policy, Corollary 5.1). All implementation details, hyperparameter configurations, and reproducibility materials are available in our public repository. [^2] https://github.com/tomatofromsky/pinn-spi-aaai2026
@@ 4.1 Linear-quadratic regulator (LQR) with compact action space
We consider entropy-regularized linear-quadratic regulator (LQR) problems in 5, 10 and 20 dimensions with compact action constraints. The system dynamics are
@@
$u\in U:=\{u\in\mathbb{R}^{m}\mid\|u\|_{\infty}\leq\textbf{u}\}.$
@@
We apply PINN-SPI and compare its performance with the model-free Soft Actor-Critic (SAC) (haarnoja2018soft) and Proximal Policy Optimization (PPO) (schulman2017proximal) algorithm. Both methods are initialized with the same linear controller derived from the unconstrained problem. Our approach uses residual minimization with randomly sampled collocation points and softmax-based policy updates.
@@
Experiments are conducted with randomly generated matrices $(A,B,Q,R)$, and set isotropic noise $\sigma=0.1I_{d}$ where $I_{d}$ denotes $d\times d$ identity matrix. We take $d=5,10$ and 20, and set $\textbf{u}=10$. Evaluation metrics include average discounted reward per iteration.
@@ Results.
Figure 1 compares PINN-SPI against SAC and PPO in 5D and 10D LQR settings. PINN-SPI consistently achieves higher reward and smoother convergence compared to other methods.
Figure 2 further illustrates the convergence behavior of PINN-SPI. The evaluation reward increases monotonically over training time, confirming the theoretical stability of policy iteration. The 10D case in particular highlights the scalability of our mesh-free implementation. An additional experiment for 20D case is provided in Appendix E.
@@
Figure 1: Comparison of PINN-SPI, SAC, and PPO on 5D and 10D stochastic LQR problems with compact action constraints.
(a) 5D stochastic LQR problem.
(b) 10D stochastic LQR problem.
@@ 4.2 Nonlinear benchmarks with stochastic dynamics
To demonstrate performance in more realistic and nonlinear settings, we evaluate our method on two standard control benchmarks: the stochastic inverted pendulum and cartpole. Each system is modeled as a stochastic control-affine system with additive Brownian noise. We adopt a standard stochasticized version of the cartpole and pendulum dynamics supported by OpenAI GYM (brockman2016openai) with additive noise on state variables.
@@
For each environment, we define entropy-regularized cost functionals and apply PINN-SPI with a fixed noise level $\sigma=0.1I_{d}$. We use neural networks for both value approximation and policy extraction, trained via policy iteration with residual loss minimization.
@@
Figure 3 shows the performance over elapsed time. PINN-SPI achieves faster stabilization and higher reward than SAC and PPO, while maintaining constraint satisfaction.
(a) Stochastic cartpole problem.
(b) Stochastic pendulum problem.
Figure 3: Comparison between PINN-SPI, SAC, and PPO on the cartpole and pendulum problems.